8z-9-2z=3=0
To get the term containing Z alone on the left side of the equation, what number needs to be added to both sides? Is it 3??

Answer by Boreal(15235)

8z-9-2z=3
6x-9=3
Add 9 to both sides
6z-9+9= 3 + 9=12
z=2